OpenStudy (anonymous):

solve: sin^2 x + sin x cos x = 0

OpenStudy (anonymous):

\[\sin ^{2} x + \sin x \cos x = 0\]

OpenStudy (anonymous):

Factoring \(\sin(x)\) out, we get: \[ \sin^2x+\sin x\cos x = (\sin(x))(\sin(x)+\cos(x))=0 \]Therefore, we have two solvable equations: \[ \sin(x)=0 \]And \[ \sin(x)=-\cos(x) \]This turns into: \[ \sin(x)=\sin(x-\frac{\pi}{2}) \]Solving each of these gives you the desired result.

OpenStudy (anonymous):

Factor out sin(x): \[\sin(x)(\sin(x)+\cos(x))=0 \implies \sin(x)=0 , \sin(x)+\cos(x)=0\]\[\sin(x)=0 \implies x = \pi k| k \in \mathbb{Z}\]\[\sin(x)+\cos(x)=0 \implies [\sin(x)+\cos(x)]^2=0 \]\[\implies \sin^2(x)+2\sin(x)\cos(x)+\cos^2(x)=0 \implies \sin(2x)+1=0\]\[\implies \sin(2x)=-1 \implies 2x =\frac{ 3 \pi }{ 2 } + 2\pi k \implies x =\frac{ 3 \pi }{ 4}+ \pi k|k \in \mathbb{Z}\] Do you understand? @jdc1980

OpenStudy (anonymous):

How did you get the sin (x-pi/2) ?

OpenStudy (anonymous):

Note that \(\sin(x)\) is an odd function and that \(\cos(x)=\sin(\frac{\pi}{2}-x)\) by the sum of angles of a right triangle. So, negating the above, we get: \[ -cos(x)=-\sin(\frac{\pi}{2}-x)=\sin(-(\frac{\pi}{2}-x))=\sin(x-\frac{\pi}{2}) \]

OpenStudy (anonymous):

Still not clicking... Is \[\csc x = \sin (\frac{ x }{ 2 } - x)\] a common identity?

OpenStudy (anonymous):

cos sorry

OpenStudy (anonymous):

It's a common identity, because the angles must add up to 90 degrees or \(\frac{\pi}{2}\), right? (Remember we're dealing with a right triangle). Hence \(\cos(x)=\sin(\frac{\pi}{2}-x\))

OpenStudy (anonymous):

okay! I see! now I'm looking for all solutions where x is 0 <= x < pi what should I do next?

OpenStudy (anonymous):

So, we have that: \[ \sin(\frac{3\pi}{4})=\sin(\frac{3\pi}{4}-\frac{\pi}{2}) \]Which is our only solution in the range \(x\in[0, \pi)\).

OpenStudy (anonymous):

This is where I'm at: